### v1

# Asymmetry in Spectral Graph Theory: Harmonic Analysis on Directed Networks via Biorthogonal Bases

Abstract
The operator-theoretic dichotomy underlying diffusion on directed networks is symmetry versus non-self-adjointness of the Markov transition operator. In the reversible (detailed-balance) regime, a directed random walk $P$ is self-adjoint in a stationary $\pi$-weighted inner product and admits orthogonal spectral coordinates; outside reversibility, $P$ is genuinely non-self-adjoint (often non-normal), and stability is governed by biorthogonal geometry and eigenvector conditioning. In this paper we develop an original harmonic-analysis framework for directed graphs anchored on the random-walk transition matrix $P=D_{\mathrm{out}}^{-1}A$ and the random-walk Laplacian $L_{\mathrm{rw}}=I-P$. Using biorthogonal left/right eigenvectors we define a Biorthogonal Graph Fourier Transform (BGFT) adapted to directed diffusion, propose a diffusion-consistent frequency ordering based on decay rates $\Re(1-\lambda)$, and derive operator-norm stability bounds for iterated diffusion and for BGFT spectral filters. We prove sampling and reconstruction theorems for $P$-bandlimited (equivalently $L_{\mathrm{rw}}$-bandlimited) signals and quantify noise amplification through the conditioning of the biorthogonal eigenbasis. A simulation protocol on directed cycles and perturbed non-normal digraphs demonstrates that asymmetry alone does not dictate instability, whereas non-normality and eigenvector ill-conditioning drive reconstruction sensitivity, making BGFT the correct analytical language for directed diffusion processes.
Key words: directed graphs; random walks; non-normal matrices; biorthogonal eigenvectors; graph Fourier transform; sampling; reversibility.
2020 Mathematics Subject Classification: Primary 05C50; Secondary 15A18, 47A10, 60J10, 94A12.
1 Introduction
1.1 Symmetry vs. non-self-adjointness: Markov operators on directed networks
A directed network naturally carries a one-step evolution operator: the random-walk (Markov) transition matrix
$P=D_{\mathrm{out}}^{-1}A,\qquad P\mathbf{1}=\mathbf{1},$
and its generator $L_{\mathrm{rw}}=I-P$. From the operator-theoretic viewpoint, the central dichotomy is not “directed vs. undirected” per se, but symmetry vs. non-self-adjointness of the Markov operator.
The symmetry regime is the reversible (detailed-balance) case: there exists a stationary distribution $\pi$ with $\Pi=\operatorname{diag}(\pi)$ such that
$\Pi P=P^{\top}\Pi,$
equivalently $P$ is self-adjoint in the weighted Hilbert space $(\mathbb{C}^{n},\langle\cdot,\cdot\rangle_{\pi})$. In that regime, $P$ is similar to a symmetric matrix $S=\Pi^{1/2}P\Pi^{-1/2}$, hence the spectrum is real and there is an orthonormal eigenbasis in the $\pi$-metric. This is precisely the mechanism by which a directed diffusion can retain symmetry (in a stationary metric), recovering Parseval-type identities and a clean variational frequency ordering. The asymmetry regime is non-reversibility, where $P$ is genuinely non-self-adjoint and may be non-normal; then orthogonality is lost, spectral coordinates can be ill-conditioned, and stability is governed by eigenvector conditioning and non-normal effects [7, 6].
Our goal is to build a harmonic-analysis calculus that is native to the Markov operator $P$: it should reduce to the classical orthogonal theory in the reversible (symmetric) regime, and it should remain exact and analyzable in the non-reversible (non-self-adjoint) regime. The correct language here is biorthogonality: left/right eigenvectors provide an exact analysis/synthesis pair even when $P$ is not normal.
1.2 Position relative to graph Fourier analysis
Graph Fourier analysis is often introduced through symmetric operators (undirected Laplacians/adjacencies), which guarantee orthogonal eigenvectors and stable spectral coordinates [2, 3, 16]. Directed graph settings typically replace symmetry by alternative constructions: optimization-based directed transforms and directed Laplacians associated with random walks and Cheeger-type theory [4, 1], as well as survey-level treatments of directed-graph signal processing [19]. Our approach is complementary and more “operator first”: we start from the canonical diffusion operator $P$ and develop an exact biorthogonal Fourier calculus for directed diffusion, with a transparent symmetry/asymmetry interpretation in terms of reversibility [9, 11].
1.3 Main contributions
Main contributions (original).
  1. 1.
    (Markov-operator BGFT) We define the Biorthogonal Graph Fourier Transform (BGFT) for the random-walk operator $P$ (equivalently $L_{\mathrm{rw}}=I-P$) via left/right eigenvectors, yielding exact analysis/synthesis identities and diagonal dynamics for diffusion iterates.
  2. 2.
    (Symmetry principle via reversibility) We identify reversibility (detailed balance) as the precise symmetry notion for directed diffusion: in the $\pi$-metric, reversible $P$ becomes self-adjoint, restoring orthogonality/Parseval identities and an exact diffusion-variational frequency ordering.
  3. 3.
    (Diffusion-consistent frequency) We propose a diffusion-consistent frequency ordering based on the decay rate $\Re(1-\lambda)$ (and magnitude alternatives), aligning with the symmetry limit and the long-time behavior of $x_{t+1}=Px_{t}$.
  4. 4.
    (Stability theorems for non-self-adjoint diffusion) We prove operator-norm bounds for diffusion iterates $P^{t}$ and for BGFT spectral filters $h(P)$, explicitly separating eigenvalue decay from eigenvector conditioning, the key instability driver in non-normal settings.
  5. 5.
    (Sampling and reconstruction) We prove sampling/reconstruction theorems for $P$-bandlimited signals and quantify noise amplification through $\sigma_{\min}(P_{M}V_{\Omega})$ and conditioning of the biorthogonal eigenbasis.
  6. 6.
    (Asymmetry vs. non-normality: numerical separation) We introduce simple indices for directedness and departure from normality and provide experiments (directed cycle vs. perturbed non-normal digraphs) showing that asymmetry alone need not cause instability, whereas non-normality and eigenvector ill-conditioning do.
1.4 Organization
Section 2 introduces directed diffusion operators and asymmetry/non-normality indices. Section 3 presents reversibility as the symmetry regime in the stationary metric. Sections 4–5 develop BGFT and stability bounds for diffusion and filtering. Section 6 proves sampling and reconstruction results, followed by algorithms and illustrative experiments.
2 Preliminaries: directed diffusion operators
2.1 Directed graphs, adjacency, and out-degree
Let $G=(V,E,w)$ be a directed weighted graph with $|V|=n$ and adjacency $A\in\mathbb{R}^{n\times n}$:
$A_{ij}=\begin{cases}w(i,j),&(i,j)\in E,\\ 0,&\text{otherwise.}\end{cases}$
Define out-degrees $d_{i}^{\mathrm{out}}=\sum_{j}A_{ij}$ and $D_{\mathrm{out}}=\operatorname{diag}(d_{1}^{\mathrm{out}},\dots,d_{n}^{\mathrm{out}})$.
2.2 Transition matrix and random-walk Laplacian
Definition 2.1 (Random-walk transition matrix).
Assume $d_{i}^{\mathrm{out}}>0$ for all $i$ (no sinks). Define
$P:=D_{\mathrm{out}}^{-1}A.$
Then $P$ is row-stochastic: $P\mathbf{1}=\mathbf{1}$.
Definition 2.2 (Random-walk Laplacian).
Define
$L_{\mathrm{rw}}:=I-P.$
Proposition 2.3 (Basic properties).
(i) $P\mathbf{1}=\mathbf{1}$ and $L_{\mathrm{rw}}\mathbf{1}=0$.  (ii) If $P$ is irreducible and aperiodic, then the diffusion $x_{t+1}=Px_{t}$ converges to the stationary component (Markov mixing perspective).
Proof.
(i) Row-stochasticity gives $P\mathbf{1}=\mathbf{1}$, hence $(I-P)\mathbf{1}=0$.
(ii) This is standard Markov chain theory; see [6, 9, 11].. ∎
2.3 Asymmetry and non-normality indices
Definition 2.4 (Asymmetry index).
For any matrix $M$, define $\alpha(M):=\left\lVert M-M^{\top}\right\rVert_{F}/\left\lVert M\right\rVert_{F}$ (with $\alpha(0)=0$).
Definition 2.5 (Departure from normality).
For any matrix $M$, define $\delta(M):=\left\lVert MM^{\ast}-M^{\ast}M\right\rVert_{F}/\left\lVert M\right\rVert_{F}^{2}$ (with $\delta(0)=0$).
Such non-normality measures (and related bounds) are classical in matrix analysis; see [12, 14, 13, 7].
We will use these for $M=P$ and $M=L_{\mathrm{rw}}$ to separate structural directedness from numerical instability drivers.
3 Reversibility as the symmetry regime for directed diffusion
Let $P\in\mathbb{R}^{n\times n}$ be row-stochastic ($P\mathbf{1}=\mathbf{1}$). Assume $P$ has a stationary distribution $\pi\in\mathbb{R}^{n}$ with $\pi_{i}>0$ and $\pi^{\top}P=\pi^{\top}$. Let $\Pi:=\mathrm{diag}(\pi)$.
Define the $\pi$-weighted inner product and norm by
$\langle x,y\rangle_{\pi}:=x^{\top}\Pi y,\qquad\|x\|_{\pi}^{2}:=\langle x,x\rangle_{\pi}.$
The adjoint of $P$ with respect to $\langle\cdot,\cdot\rangle_{\pi}$ is
$P^{\dagger}:=\Pi^{-1}P^{\top}\Pi,\quad\text{so that}\quad\langle Px,y\rangle_{\pi}=\langle x,P^{\dagger}y\rangle_{\pi}.$
Definition 3.1 (Reversibility / detailed balance).
$P$ is reversible (w.r.t. $\pi$) if
$\Pi P=P^{\top}\Pi,\quad\text{equivalently}\quad P=P^{\dagger}.$
This detailed-balance condition is standard in reversible Markov chain theory; see [10, 9, 6].
Theorem 3.2 (Weighted symmetry equivalences).
The following are equivalent:
  1. (i)
    $P$ is reversible: $\Pi P=P^{\top}\Pi$.
  2. (ii)
    $P$ is self-adjoint in $\langle\cdot,\cdot\rangle_{\pi}$: $P=P^{\dagger}$.
  3. (iii)
    The similarity transform $S:=\Pi^{1/2}P\Pi^{-1/2}$ is symmetric: $S=S^{\top}$.
In this case, $P$ has a complete $\pi$-orthonormal eigenbasis, and all eigenvalues are real.
Proof.
[…]
See also [9, 11] for related equivalences and consequences.
Remark 3.3 (Symmetry/asymmetry interpretation for this paper).
Undirected diffusion is symmetric in the standard Euclidean inner product. Directed diffusion can still be symmetric in the weighted $\pi$-inner product exactly in the reversible regime. Non-reversibility is the correct notion of asymmetry for random-walk harmonic analysis.
4 Biorthogonal Graph Fourier Transform (BGFT) for random-walk diffusion
4.1 Left/right eigenvectors and BGFT
Assumption 4.1 (Diagonalizability).
Assume $P$ is diagonalizable over $\mathbb{C}$:
$P=V\Lambda V^{-1},\qquad\Lambda=\operatorname{diag}(\lambda_{1},\dots,\lambda_{n}),$
with right eigenvectors $V=[v_{1}\ \cdots\ v_{n}]$.
Define $U^{\ast}:=V^{-1}$ so $U^{\ast}V=I$ and $u_{k}^{\ast}v_{\ell}=\delta_{k\ell}$.
Definition 4.2 (BGFT (diffusion version)).
For a graph signal $x\in\mathbb{C}^{n}$, define BGFT coefficients
$\widehat{x}:=U^{\ast}x,\qquad\widehat{x}_{k}=u_{k}^{\ast}x,$ (1)
and synthesis
$x=V\widehat{x}=\sum_{k=1}^{n}v_{k}\,\widehat{x}_{k}.$ (2)
Theorem 4.3 (Perfect reconstruction).
Under Assumption 4.1, for all $x\in\mathbb{C}^{n}$,
$I=\sum_{k=1}^{n}v_{k}u_{k}^{\ast},\qquad x=\sum_{k=1}^{n}v_{k}\,u_{k}^{\ast}x.$
Proof.
Since $U^{\ast}V=I$, we have $VU^{\ast}=I$; expand $VU^{\ast}$ in columns/rows. ∎
4.2 Diffusion dynamics are diagonal in BGFT coordinates
Theorem 4.4 (BGFT-domain diffusion).
Let $x_{t+1}=Px_{t}$ with $x_{0}\in\mathbb{C}^{n}$. Then
$\widehat{x}_{t}=U^{\ast}x_{t}=\Lambda^{t}\,\widehat{x}_{0},\qquad x_{t}=V\Lambda^{t}U^{\ast}x_{0}.$
Equivalently, for $L_{\mathrm{rw}}=I-P$,
$\widehat{(L_{\mathrm{rw}}x)}=(I-\Lambda)\widehat{x}$.
Proof.
Use $P=V\Lambda U^{\ast}$ and $U^{\ast}V=I$. Then $U^{\ast}(Px)=\Lambda(U^{\ast}x)$ and iterate. ∎
4.3 Diffusion-consistent frequency ordering
For diffusion, the mode with eigenvalue $\lambda$ evolves as $\lambda^{t}$. If $|\lambda|<1$, it decays; if $\lambda\approx 1$, it is slowly varying (low frequency). We define the diffusion decay rate:
$\omega_{\mathrm{diff}}(\lambda):=\Re(1-\lambda).$
Low $\omega_{\mathrm{diff}}$ corresponds to persistent/slow modes; high $\omega_{\mathrm{diff}}$ corresponds to fast decay. In the symmetric undirected limit (where $\lambda\in[-1,1]$ for normalized settings), this ordering aligns with classical low/high frequency intuition.
5 Directed diffusion filtering and stability bounds
5.1 Spectral filters
Definition 5.1 (BGFT spectral filter for diffusion).
Let $h:\mathbb{C}\to\mathbb{C}$. Define
$H:=V\,h(\Lambda)\,U^{\ast}.$
Proposition 5.2 (Diagonal action in BGFT domain).
For $\widehat{x}=U^{\ast}x$,
$\widehat{Hx}=U^{\ast}Hx=h(\Lambda)\widehat{x}$.
Proof.
Compute $U^{\ast}Vh(\Lambda)U^{\ast}x=h(\Lambda)\widehat{x}$. ∎
5.2 Operator-norm stability: diffusion and filtering
Theorem 5.3 (Norm bound for diffusion iterates).
Assume $P=V\Lambda V^{-1}$. Then for every $t\in\mathbb{N}$,
$\left\lVert P^{t}\right\rVert_{2}\leq\operatorname{cond}(V)\,\max_{k}|\lambda_{k}|^{t},\qquad\operatorname{cond}(V)=\left\lVert V\right\rVert_{2}\left\lVert V^{-1}\right\rVert_{2}.$
Proof.
$P^{t}=V\Lambda^{t}V^{-1}$, hence $\left\lVert P^{t}\right\rVert_{2}\leq\left\lVert V\right\rVert_{2}\left\lVert\Lambda^{t}\right\rVert_{2}\left\lVert V^{-1}\right\rVert_{2}=\operatorname{cond}(V)\max_{k}|\lambda_{k}|^{t}$. ∎
Theorem 5.4 (Norm bound for spectral filters).
Let $H=Vh(\Lambda)V^{-1}$. Then
$\left\lVert H\right\rVert_{2}\leq\operatorname{cond}(V)\,\max_{k}|h(\lambda_{k})|.$
Proof.
$\left\lVert H\right\rVert_{2}\leq\left\lVert V\right\rVert_{2}\,\left\lVert h(\Lambda)\right\rVert_{2}\,\left\lVert V^{-1}\right\rVert_{2}=\operatorname{cond}(V)\max_{k}|h(\lambda_{k})|$. ∎
Remark 5.5 (Symmetry/asymmetry interpretation).
When $P$ is normal and diagonalizable by a unitary basis, $\operatorname{cond}(V)=1$ and the bounds become tight and symmetry-like. For non-normal $P$, $\operatorname{cond}(V)$ can be large, creating instability even if $|\lambda_{k}|\leq 1$.Sharper growth control for non-normal matrices can be expressed via pseudospectral/Kreiss-type constants; see [15, 7].
6 BGFT energy in the stationary metric and its symmetry limit
Assume $P$ is diagonalizable over $\mathbb{C}$: $P=V\Lambda V^{-1}$ and define $U^{*}:=V^{-1}$. Let $\widehat{x}:=U^{*}x$ be BGFT coefficients so that $x=V\widehat{x}$.
Theorem 6.1 ($\pi$-metric Parseval identity).
For any $x\in\mathbb{C}^{n}$,
$\|x\|_{\pi}^{2}=\widehat{x}^{\,*}\,G_{\pi}\,\widehat{x},\qquad G_{\pi}:=V^{*}\Pi V.$
Proof.
Since $x=V\widehat{x}$, $\|x\|_{\pi}^{2}=x^{*}\Pi x=\widehat{x}^{*}(V^{*}\Pi V)\widehat{x}$. ∎
Corollary 6.2 (Two-sided bounds via conditioning in $\pi$).
Let $W:=\Pi^{1/2}V$. Then
$\sigma_{\min}(W)^{2}\|\widehat{x}\|_{2}^{2}\leq\|x\|_{\pi}^{2}\leq\sigma_{\max}(W)^{2}\|\widehat{x}\|_{2}^{2}.$
Equivalently, energy distortion is controlled by $\kappa(W)=\sigma_{\max}(W)/\sigma_{\min}(W)$.
6.1 Diffusion variation and frequency ordering
Define the random-walk Laplacian $L_{\mathrm{rw}}:=I-P$ and the diffusion variation
$\mathrm{TV}_{\pi}(x):=\|L_{\mathrm{rw}}x\|_{\pi}^{2}=\|(I-P)x\|_{\pi}^{2}.$
Theorem 6.3 (BGFT-domain bounds for diffusion variation).
With $x=V\widehat{x}$ and $W=\Pi^{1/2}V$,
$\sigma_{\min}(W)^{2}\sum_{k=1}^{n}|1-\lambda_{k}|^{2}|\widehat{x}_{k}|^{2}\;\leq\;\mathrm{TV}_{\pi}(x)\;\leq\;\sigma_{\max}(W)^{2}\sum_{k=1}^{n}|1-\lambda_{k}|^{2}|\widehat{x}_{k}|^{2}.$
Proof.
$(I-P)x=V(I-\Lambda)\widehat{x}$. Then $\|(I-P)x\|_{\pi}=\|\Pi^{1/2}V(I-\Lambda)\widehat{x}\|_{2}=\|W(I-\Lambda)\widehat{x}\|_{2}$. Apply $\sigma_{\min}(W)\|z\|_{2}\leq\|Wz\|_{2}\leq\sigma_{\max}(W)\|z\|_{2}$ to $z=(I-\Lambda)\widehat{x}$ and square. ∎
Remark 6.4 (Exact symmetry limit).
If $P$ is reversible, one can choose $V$ $\pi$-orthonormal, hence $W=\Pi^{1/2}V$ is unitary and $\sigma_{\min}(W)=\sigma_{\max}(W)=1$. Then the inequalities become equalities and $|1-\lambda_{k}|$ becomes an exact diffusion frequency.
7 Sampling and reconstruction for diffusion-bandlimited signals
Let $\Omega\subset\{1,\dots,n\}$ with $|\Omega|=K$ represent the “low diffusion-frequency” modes (e.g. smallest $\omega_{\mathrm{diff}}(\lambda)$ or largest $\Re(\lambda)$). Let $V_{\Omega}\in\mathbb{C}^{n\times K}$ contain $\{v_{k}\}_{k\in\Omega}$.
Definition 7.1 (Diffusion-bandlimited signals).
A signal $x$ is $\Omega$-bandlimited (relative to $P$) if $x=V_{\Omega}c$ for some $c\in\mathbb{C}^{K}$.
Bandlimited sampling on graphs has a substantial literature; see, e.g., [18, 17].
Let $M\subset V$, $|M|=m$, and $P_{M}\in\{0,1\}^{m\times n}$ be the restriction operator.
Theorem 7.2 (Exact recovery).
If $x=V_{\Omega}c$ and $P_{M}V_{\Omega}$ has full column rank $K$, then $x$ is uniquely determined by samples $y=P_{M}x$ and recovered by
$\widehat{c}=(P_{M}V_{\Omega})^{\dagger}y,\qquad\widehat{x}=V_{\Omega}\widehat{c}$.
Proof.
Full column rank makes $P_{M}V_{\Omega}$ injective; solve the linear system in least squares. ∎
Related sampling-set conditions and reconstruction stability on graphs are discussed in [18, 17].
Theorem 7.3 (Noise sensitivity).
If $y=P_{M}x+\eta$, then the least-squares reconstruction satisfies
$\left\lVert\widehat{x}-x\right\rVert_{2}\leq\left\lVert V_{\Omega}\right\rVert_{2}\,\left\lVert(P_{M}V_{\Omega})^{\dagger}\right\rVert_{2}\,\left\lVert\eta\right\rVert_{2}=\left\lVert V_{\Omega}\right\rVert_{2}\,\frac{\left\lVert\eta\right\rVert_{2}}{\sigma_{\min}(P_{M}V_{\Omega})}.$
Proof.
$\widehat{c}-c=(P_{M}V_{\Omega})^{\dagger}\eta$ and $\widehat{x}-x=V_{\Omega}(\widehat{c}-c)$. ∎
8 Algorithms
Algorithm 1 BGFT for random-walk diffusion
1:AA (directed adjacency), $D_{\mathrm{out}}$ invertible, signal $x$
2:BGFT coefficients $\widehat{x}$, eigenpairs $(\Lambda,V)$
[…]
4:Compute eigendecomposition $P=V\Lambda V^{-1}$ (complex arithmetic)
[…]
7:return $(\widehat{x},\Lambda,V)$
Algorithm 2 Diffusion filtering and bandlimited reconstruction
[…]
2:filtered signal HxHx or reconstructed $\widehat{x}$
3:Filtering: $H\leftarrow Vh(\Lambda)V^{-1}$, output $Hx\leftarrow Hx$
4:Reconstruction: form $V_{\Omega}$, solve $\widehat{c}\leftarrow\arg\min_{c}\left\lVert P_{M}V_{\Omega}c-y\right\rVert_{2}^{2}$
[…]
9 Experiments: directed cycle vs perturbed non-normal digraphs
9.1 Graphs
Use $n\in\{32,64,128\}$ and compare:
  1. 1.
    Undirected cycle $C_{n}$ (convert to diffusion by symmetrizing and normalizing).
  2. 2.
    Directed cycle $\overrightarrow{C_{n}}$: $P$ is a permutation shift (asymmetric but normal/unitary).
  3. 3.
    Perturbed directed cycle $\overrightarrow{C_{n}}^{(\varepsilon)}$: add a directed chord then renormalize rows to keep $P$ stochastic; this typically yields non-normal $P$ and large $\operatorname{cond}(V)$.
9.2 Tasks and metrics
  • Diffusion filtering: low-pass via $h(\lambda)=\exp(-\tau(1-\lambda))$ (BGFT-defined), compare smoothing strength on the three graphs.
  • Forecasting: iterate diffusion $x_{t+1}=Px_{t}$ and compare $\left\lVert x_{t}\right\rVert_{2}$ trends with Theorem 5.3.
  • Sampling/reconstruction: generate $\Omega$-bandlimited signals and recover from $m$ samples; report RelErr and $\sigma_{\min}(P_{M}V_{\Omega})$.
Report tables/figures:
$\alpha(P),\ \delta(P),\ \operatorname{cond}(V),\ \operatorname{cond}(P_{M}V_{\Omega}),\ \mathrm{RelErr}=\frac{\left\lVert\widehat{x}-x\right\rVert_{2}}{\left\lVert x\right\rVert_{2}}.$
Table 1: Minimal numerical illustration for the transition-operator BGFT.
Graph $\alpha(P)$ $\delta(P)$ $\kappa(V)$ $\kappa(P_{M}V_{\Omega})$ RelErr
Undirected cycle $A_{\mathrm{und}}$ 0 0 1.2453204511204569 36.59492056037454 0.0000031215500108761767
Directed cycle $A_{\rightarrow}$ 1.4142135623730951 0 1 93.04681515171424 0.000007080903224516353
Perturbed $A_{\varepsilon}$ ($\varepsilon=20$) 1.414213562373095 0.02987165083714049 28.011585066632986 352.8935063092261 0.00002523914083929862
Observed separation. The directed cycle is asymmetric ($\alpha(P)>0$) but normal ($\delta(P)=0$) with a well-conditioned eigenbasis ($\kappa(V)=1$), whereas the perturbed digraph remains asymmetric but becomes non-normal ($\delta(P)>0$) and strongly ill-conditioned ($\kappa(V)\gg 1$), leading to markedly larger reconstruction error, consistent with the stability bounds.
10 Conclusion
We developed an original diffusion-centered harmonic analysis for directed graphs using the random-walk transition matrix $P$ and Laplacian $L_{\mathrm{rw}}=I-P$. The BGFT provides exact analysis/synthesis, diagonalizes diffusion dynamics, motivates a diffusion-consistent frequency ordering, and yields explicit stability bounds for iterated diffusion and spectral filtering governed by eigenvector conditioning. Sampling and reconstruction theorems quantify how non-normality amplifies noise through $\sigma_{\min}(P_{M}V_{\Omega})$ and $\operatorname{cond}(V)$. This establishes a principled symmetry/asymmetry narrative: symmetry yields orthogonality and stability; asymmetry forces biorthogonal geometry; non-normality determines practical robustness.
Acknowledgements
The author expresses his gratitude to the Commissioner of Collegiate Education (CCE), Government of Andhra Pradesh, and the Principal, Government College (Autonomous), Rajahmundry, for continued support and encouragement.
Author Contributions
The author is solely responsible for conceptualization, methodology, analysis, software, validation, and writing.
Funding
No external funding was received.
Data Availability Statement
No external datasets were used. The code that generates the reported numerical table/figures will be provided as a reproducible script and (upon acceptance) via a public repository link.
Conflicts of Interest
The author declares no conflicts of interest.
Appendix A Reference Python code (P and Lrw, BGFT, filtering, reconstruction)
import numpy as np
def directed_cycle_A(n):
A = np.zeros((n,n), dtype=float)
for i in range(n):
A[i, (i+1)%n] = 1.0
return A
def undirected_cycle_A(n):
A = np.zeros((n,n), dtype=float)
for i in range(n):
A[i, (i+1)%n] = 1.0
A[(i+1)%n, i] = 1.0
return A
def add_directed_chord(A, eps=0.2, i=0, j=None):
n = A.shape[0]
if j is None:
j = n//2
B = A.copy()
B[i, j] += eps
return B
def D_out(A):
d = A.sum(axis=1)
if np.any(d == 0):
raise ValueError("Foundsinknode(out-degree0).Fixbyaddingsmalloutgoingweight.")
return np.diag(d)
def transition_P(A):
D = D_out(A)
return np.linalg.inv(D) @ A
def L_rw(P):
n = P.shape[0]
return np.eye(n) - P
def asymmetry_index(M):
den = np.linalg.norm(M, ord=’fro’)
if den == 0:
return 0.0
return np.linalg.norm(M - M.T, ord=’fro’) / den
def departure_from_normality(M):
Mf = np.linalg.norm(M, ord=’fro’)
if Mf == 0:
return 0.0
MMstar = M @ M.conj().T
MstarM = M.conj().T @ M
return np.linalg.norm(MMstar - MstarM, ord=’fro’) / (Mf**2)
def bgft_decomposition(P):
lam, V = np.linalg.eig(P)
Vinv = np.linalg.inv(V)
Ustar = Vinv
return lam, V, Ustar
def diffusion_filter_matrix(P, tau=2.0):
# h(lambda) = exp(-tau*(1-lambda))
lam, V, Ustar = bgft_decomposition(P)
h = np.exp(-tau*(1.0 - lam))
H = V @ np.diag(h) @ Ustar
return H
def sample_operator(n, M):
m = len(M)
Pm = np.zeros((m,n), dtype=float)
for r, idx in enumerate(M):
Pm[r, idx] = 1.0
return Pm
def reconstruct_bandlimited(P, Omega, M, x, noise=0.0, seed=0):
np.random.seed(seed)
lam, V, _ = bgft_decomposition(P)
V_O = V[:, Omega]
Pm = sample_operator(P.shape[0], M)
y = Pm @ x
if noise > 0:
y = y + noise * np.random.randn(*y.shape)
B = Pm @ V_O
c_hat, *_ = np.linalg.lstsq(B, y, rcond=None)
x_hat = V_O @ c_hat
relerr = np.linalg.norm(x_hat - x) / np.linalg.norm(x)
condB = np.linalg.cond(B)
return x_hat, relerr, condB
if __name__ == "__main__":
n = 64
A_und = undirected_cycle_A(n)
A_dir = directed_cycle_A(n)
A_per = add_directed_chord(A_dir, eps=20, i=0, j=n//2)
for name, A in [("undirected", A_und), ("directed", A_dir), ("perturbed", A_per)]:
P = transition_P(A)
lam, V, _ = bgft_decomposition(P)
print(name,
"alpha(P)=", asymmetry_index(P),
"delta(P)=", departure_from_normality(P),
"cond(V)=", np.linalg.cond(V),
"rho(P)~=", np.max(np.abs(lam)))
# Create a bandlimited signal on perturbed digraph
P = transition_P(A_per)
lam, V, _ = bgft_decomposition(P)
K = 8
# Choose "low diffusion-frequency": largest Re(lambda) (closest to 1)
Omega = np.argsort(-np.real(lam))[:K]
c = np.random.randn(K) + 1j*np.random.randn(K)
x = V[:, Omega] @ c
m = 20
M = np.sort(np.random.choice(n, size=m, replace=False))
x_hat, relerr, condB = reconstruct_bandlimited(P, Omega, M, x, noise=0.0)
print("RelErr=", relerr, "cond(P_MV_Omega)=", condB)
# Diffusion smoothing example
H = diffusion_filter_matrix(P, tau=2.0)
x_smooth = H @ x
print("||x||2=", np.linalg.norm(x), "||Hx||2=", np.linalg.norm(x_smooth))